In how many ways a committee of 3 students can be selected from 4 students A,B,C,D ?
1
Expert's answer
2020-04-24T19:08:10-0400

The total number of ways=

"{C_3}^4=\frac{4!}{3!×(4-3)!}=\frac{4×3!}{3!×1!}=\frac{4}{1!}=4"